OpenStudy (anonymous):

-3/4c+1/2=3/2c-1 help!!!!

OpenStudy (anonymous):

c=3/2

OpenStudy (anonymous):

Thats what I got but mathway says its 2/3

OpenStudy (anonymous):

nope is 3/2 dun worry that correct 2/3 is wrong....

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

no problem....haha :)

OpenStudy (anonymous):

-3/4c+1/2=3/2c-1 -6/4c+1=3/2c-1 -12c+6=12c+3 -24c=-3 c=1/8, guys,is this wrong??